Investors require a 6% rate of return on Levine Company's stock (i.e., rs = 6%).

  1. What is its value if the previous dividend was D0 = $2.50 and investors expect dividends to grow at a constant annual rate of (1) -3%, (2) 0%, (3) 4%, or (4) 5%? Do not round intermediate calculations. Round your answers to the nearest cent.

    1. $ =

    2. $=

    3. $=

    4. $=
Answer #1

Current price=D1/(Required return-Growth rate)

a.Current price=(2.5*(1-0.03)/(0.06-(0.03))

=(2.425/0.09)=$26.94(Approx).

b.Current price=2.5/0.06=$41.67(Approx).

c.Current price=(2.5*1.04)/(0.06-0.04)=$130

d.Current price=(2.5*1.05)/(0.06-0.05)=$262.5